# Cosmic Ray Magnetohydrodynamics: A New Two-Moment Framework with Numerical Implementation

### v1

@@ 2 CR hydrodynamics
2.1 CR Evolution in Phase Space
@@
where $f\equiv f(\vec{x},\vec{p},t)$ is the CR distribution and $\vec{v}$ is the CR particle velocity. The net force $d\vec{p}/dt$ includes the dominant Lorentz force and a remaining force due to the Alfvén waves which is treated as a scattering process. In a frame comoving with the MHD gas, we benefit from the fact that CRs’ gyromotion is axisymmetric about field lines, which is in general not true in the lab frame where CRs are transported across field lines. For this reason, we transform equation Equation 2 into a frame that moves at $\vec{u}_{\perp}\equiv\vec{u}-\vec{u}\cdot\vec{b}\vec{b}$, where $\vec{u}$ is the background gas velocity and $\vec{b}\equiv\vec{B}/B$ is the unit vector along magnetic field. This frame is as close to the lab frame as possible while preserving the gyrotropic symmetry, because in the direction perpendicular to the magnetic field, this frame has no relative motion with respect to the field lines.
In this reference frame, the CR distribution has no dependence on the gyrophase $\phi$, and we can write it as $\bar{f}(\vec{x},p,\mu,t)$ where $p$ is the magnitude of particle momentum and $\mu$ is the pith-angle cosine $(\mu\equiv\vec{p}\cdot\vec{b}/p)$. After averaging over $\phi$, Skilling (1975) writes the evolution equation in the frame moving at $\vec{u}_{\perp}$ as his equation (5). We find this is equivalent to the following form [^1] This result is consistent with equation (5.21) in Zank (2014) and equation (29) in Thomas & Pfrommer (2019), except they ignored the $\partial\vec{b}/\partial t$ terms:
@@
By keeping only the dominant terms in equation Equation 5, we can rewrite the scattering term into the Fokker-Plank form [^2] Expanding Equation 5 in fact gives $\sum_{i=\pm}\partial\left[\left(1-3\mu\vec{b}\cdot\vec{w}_{\pm}/c\right)F_{\mu,i}\right]/\partial\mu+{p^{-2}}\partial(p^{2}F_{p,i})/\partial p$. However, when we take moments, the $3\mu\vec{b}\cdot\vec{w}_{\pm}/c$ term is always much less than order unity and we can safely omit that term.:
@@
2.2 CR Fluid Dynamics
In this subsection we first introduce some definitions of our method, and then formulate the CR fluid equations by taking the traditional moment hierarchy. From here we omit the spatial coordinate in the CR distribution function since we will mostly focus on integrals over momentum space which are spatially local.
@@ 2.2 CR Fluid Dynamics
Even if the level of anisotropy is quite small, $\Delta P_{\rm cr}$ can play an important role in both macro- and microphysical processes. Microscopically, the CR pressure anisotropy can trigger gyroresonant instability (Kulsrud, 2004; Yan & Lazarian, 2011; Lazarian & Beresnyak, 2006; Sun et al., 2024) similar to the CRSI driven by a drift anisotropy, and thus bring in a different type of self-confinement. This mechanism can potentially have many macroscopic consequences. For example in the case of CR bottlenecks (e.g, Skilling, 1971; Wiener et al., 2017) where the $\mathcal{E}_{\rm cr}$ gradient is flattened and the streaming instability is deactivated, the CRPAI may then dominate and govern CR momentum transfer to the gas, which may alter the properties of CR-driven wind. In situations such as a supernova remnant where the system undergoes rapid expansion, the CR pressure anisotropy could be efficiently driven and induce a high level of heating (Zweibel, 2020). Additionally, consider a magnetic field with a focusing structure and suppose there is no MHD wave scattering at all, then CR flux should still be affected by the $\Delta P_{\rm cr}$ in the advection term $\nabla\cdot\mathbf{P}_{\rm cr}$ due to adiabatic invariance.
@@
Expanding CR distributions $f$ by Legendre polynomials is canonical in the study about CR transport because of their natural geometric characteristics in pitch angle (e.g., Klimas & Sandri, 1971; Earl, 1973; Webb, 1987; Zank et al., 2000; Snodin et al., 2006; Litvinenko & Noble, 2013; Rodrigues et al., 2019; Thomas & Pfrommer, 2019; Zweibel, 2020). Mathematically, a rigorous expansion certainly must involve a complete set of Legendre polynomials as basis functions. However, from the definitions in Equation 11, Equation 14, Equation 18 and the relation ${\rm tr}(\mathbf{P}_{\rm cr})=\mathcal{E}_{\rm cr}$, we see Legendre polynomials up to the second order are already sufficient to incorporate $\Delta P_{\rm cr}$. We thus truncate at second order. Given the orthogonality of Legendre polynomials, each $f_{n}(p,t)\ (n=0,1,2)$ component separately corresponds to a fluid moment:
@@
With these conventions, we now take the $\mathcal{E}_{\rm cr}$ and $F_{\rm cr,\parallel}/c^{2}$ moments on top of the kinetic equation Equation 3, yielding left hand sides (LHSs) as:
@@ 2.3 CR Scattering
The values of these terms depend on the intensities of the corresponding resonant waves (Schlickeiser, 1989), which are determined by specific physical processes. However, even without a specified wave-driving mechanism, from the resonant condition $k\mu v\approx\pm\Omega_{\rm cr}/\gamma$ we can see forward/backward propagating particles $(\mu>0/\mu<0)$ only resonantly interact with right-hand/left-hand polarized waves $(+/-\ {\rm in\ front\ of\ }\Omega_{\rm cr})$, such that:
@@
The exact value of the factor $\alpha$ can be calibrated from MHD-PIC simulations (e.g., Bai, 2022; Sun & Bai, 2023).
@@
$+5\alpha v_{A}c[(\sigma_{+}^{L}-\sigma_{+}^{R})w_{+,\parallel}-(\sigma_{-}^{L}-\sigma_{-}^{R})w_{-,\parallel}]\Delta P_{\rm cr}$
@@
$+5\alpha\left(\sigma_{+}^{L}-\sigma_{+}^{R}+\sigma_{-}^{L}-\sigma_{-}^{R}\right)c\Delta P_{\rm cr}$
@@
2.4 CR Hydrodynamic Equations in Lab Frame
@@
As CRs are relativistic particles, CR hydrodynamics shares similarities to radiation transport particularly in frame transformation relations. Thomas & Pfrommer (2019) showed the results for CRs by combining $\mathcal{E}_{\rm cr}$, $\vec{F}_{\rm cr}$ and $\mathbf{P}_{\rm cr}$ as a rank-2 energy momentum tensor and then applying a Lorentz transformation. We slightly extended their results to the case with an anisotropic $\mathbf{P}_{\rm cr}$, and list the relations below. Our calculation and final results are in direct analogy with the relations in radiation transport (e.g., Mihalas & Mihalas, 1984; Castor, 2007; Jiang et al., 2012):
@@
One issue associated with transforming $\mathbf{P}_{\rm cr}$ to the lab frame is the difficulty of determining $\Delta P_{\rm cr}$ in the lab frame. In fact it makes even no sense to define $\Delta P_{\rm cr}$ in the lab frame because the CR distribution $f$ is no longer symmetric about the gyro phase. However, quasi-linear theory suggests that the anisotropic component in the ansatz Equation 19 will approximately satisfy $|f_{1}/f_{0}|\sim v_{A}/c$ and $|f_{2}/f_{0}|\sim v_{A}/c$ (Sun et al., 2024), which by equations Equation 20-Equation 22 implies $|\vec{F}_{\rm cr}|\sim v_{A}\mathcal{E}_{\rm cr}$ and $|\Delta P_{\rm cr}|\sim v_{A}\mathcal{E}_{\rm cr}/c$ (see also Section Section 2.5). Consequently, the additional term arising from $\mathbf{P}_{\rm cr}$ transformation is $\sim\mathcal{O}(v_{A}u_{\perp}/c^{2})$, which is generally significantly smaller than $\Delta P_{\rm cr}$. We can thus consider $\mathbf{P}_{\rm cr,lab}\approx\mathbf{P}_{\rm cr}$, and regard $\Delta P_{\rm cr}$ somewhat as a scalar. These considerations imply that the closure for $\mathbf{P}_{\rm cr,lab}$ becomes
@@
Thomas & Pfrommer (2019) suggested this Lorentz force should counteract the perpendicular CR inertia and establish a dynamical equilibrium $\vec{j}_{\rm cr}\times\vec{B}/c=\nabla_{\perp}P_{\rm cr}$ on a kinetic time scale, which keeps the $\vec{F}_{\rm cr}$ in the comoving frame parallel to the magnetic field. Given such a consideration and the fact that Equation 24 is obviously the parallel component of $c^{-2}\partial\vec{F}_{\rm cr,lab}/\partial t+\nabla\cdot\mathbf{P}_{\rm cr,lab}$, we expect the additional perpendicular component to balance the Lorentz force, and thus write down the full CR flux equation as:
@@
$+\left[(\sigma_{+}^{L}-\sigma_{-}^{R})(u_{\parallel}+v_{A})+(\sigma_{-}^{L}-\sigma_{-}^{R})(u_{\parallel}-v_{A})\right]c\Delta P_{\rm cr}$
@@ 2.5 Subgrid Physics
    A large-scale CR pressure gradient $\nabla P_{\rm cr}$ drives CR bulk motion (streaming), leading to the CRSI when this bulk motion exceeds the Alfvén speed. The CRSI excites two branches of the Alfvén waves propagating down the direction of $\nabla P_{\rm cr}$ projected along the magnetic field (Kulsrud & Pearce, 1969; Wentzel, 1974). Given a wave damping rate $\nu_{\rm{damp}}$, the scattering coefficients have been shown to be (Kulsrud, 2004; Bai et al., 2019; Bai, 2022; Armillotta et al., 2021):
@@
    where $\nu_{\rm damp}$ is the damping rate and $\rho$ is the gas density.
    With this prescription, the resulting equations are in alignment with the current two-moment CR fluid formulation (e.g. Jiang & Oh, 2018; Thomas & Pfrommer, 2019) as expected in the situation solely subject to the CRSI.
    The above relation applies for either linear or nonlinear wave damping (e.g. ion-neutral or nonlinear Landau; (Kulsrud & Pearce, 1969)). For ion-neutral damping, appropriate in neutral gas, the resulting $\sigma_{\pm}^{L,R}$ is linear in $|\nabla P_{\rm cr}|$; for nonlinear Landau damping, appropriate in ionized gas, $\sigma_{\pm}^{L,R}$ varies as $\sqrt{|\nabla P_{\rm cr}|}$ (Armillotta et al., 2021).
@@
    The CRPAI is a less studied version of the CR gyro-resonant instability, which can be triggered when the level of CR pressure anisotropy $\Delta P_{\rm cr}/P_{\rm cr}\gtrsim v_{A}/c$ (Kulsrud, 2004; Zweibel, 2020). It triggers one branch of Alfvén wave along each propagating direction. In reality, this may occur when the bulk CRs experience a varying magnetic field, e.g., when CRs travel through magnetic fields with large spatial variations, or the background field undergoes shear motion. Recent work (e.g., Sun et al., 2024) has isolated this effect using an expanding/compressing box setup. They derived the scattering coefficients based on quasi-linear theory, and calibrated the results using MHD-PIC simulations. The result is:
@@
The other ingredient for user input is the value of CR pressure anisotropy $\Delta P_{\rm cr}$, which is again associated with the CRPAI. Similarly based on quasi-linear theory, by balancing wave growth and (linear) damping, together with the growth of pressure anisotropy and quasi-linear diffusion, the steady state value of $\Delta P_{\rm cr}$ can be estimated as (Sun et al., 2024):
@@
The +/+/- signs correspond to a decreasing/increasing magnetic field strength that makes $P_{\rm cr,\perp}$ smaller/larger than the $P_{\rm cr,\parallel}$.
We highlight two important points behind the aforementioned prescriptions. First, the scattering coefficients and pressure anisotropy are derived assuming steady-state balance relations between driving and wave damping. Typically, this balance can be established on an instability growth timescale that is much shorter than the dynamic timescale over the macroscopic system. Therefore, it usually suffices to adopt such steady-state estimates. Second, we note that these prescriptions were calculated by assuming either the CRSI or CRPAI being the only instability at work. In reality, the CRPAI and CRSI are likely to act together, although the outcome is not additive, because each of them amplify (different) two out of four parallel-propagating Alfvén waves, while damping the other two. For present purposes, we only consider the case with pure CRSI or CRPAI, while leaving the more general case to future work.
3 Numerical Implementation

### v2

@@ II CR hydrodynamics
II.1 CR Evolution in Phase Space
@@
where $f\equiv f(\vec{x},\vec{p},t)$ is the CR distribution and $\vec{v}$ is the CR particle velocity. The net force $d\vec{p}/dt$ includes the dominant Lorentz force and a remaining force due to the Alfvén waves which is treated as a scattering process. In a frame comoving with the MHD gas, we benefit from the fact that CRs’ gyromotion is axisymmetric about field lines, which is in general not true in the lab frame where CRs are transported across field lines. For this reason, we transform Equation 2 into a frame that moves at $\vec{u}_{\perp}\equiv\vec{u}-\vec{u}\cdot\vec{b}\vec{b}$, where $\vec{u}$ is the background gas velocity and $\vec{b}\equiv\vec{B}/B$ is the unit vector along magnetic field. This frame is as close to the lab frame as possible while preserving the gyrotropic symmetry, because in the direction perpendicular to the magnetic field, this frame has no relative motion with respect to the field lines.
In this reference frame, the CR distribution has no dependence on the gyrophase $\phi$, and we can write it as $\bar{f}(\vec{x},p,\mu,t)$ where $p$ is the magnitude of particle momentum and $\mu$ is the pitch-angle cosine $(\mu\equiv\vec{p}\cdot\vec{b}/p)$. After averaging over $\phi$, Skilling (1975) writes the evolution equation in the frame moving at $\vec{u}_{\perp}$ as his equation (5). We find this is equivalent to the following form [^1] This result is consistent with equation (5.21) in Zank (2014) and equation (29) in Thomas and Pfrommer (2019), except they ignored the $\partial\vec{b}/\partial t$ terms:
@@
By keeping only the dominant terms in Equation 5, we can rewrite the scattering term into the Fokker-Plank form [^2] Expanding Equation 5 in fact gives $\sum_{i=\pm}\partial\left[\left(1-3\mu\vec{b}\cdot\vec{w}_{\pm}/c\right)F_{\mu,i}\right]/\partial\mu+{p^{-2}}\partial(p^{2}F_{p,i})/\partial p$. However, when we take moments, the $3\mu\vec{b}\cdot\vec{w}_{\pm}/c$ term is always much less than order unity and we can safely omit that term.:
@@
II.2 CR Fluid Dynamics
In this subsection we first introduce some definitions of our method, then formulate the CR fluid equations by taking the traditional moment hierarchy. From here we omit the spatial coordinate in the CR distribution function since we will mostly focus on integrals over momentum space which are spatially local.
@@ II.2 CR Fluid Dynamics
Even if the level of anisotropy is quite small, $\Delta P_{\rm cr}$ can play an important role in both macro- and microphysical processes. Microscopically, the CR pressure anisotropy can trigger gyroresonant instability (Kulsrud, 2004; Yan and Lazarian, 2011; Lazarian and Beresnyak, 2006; Sun et al., 2024) similar to the CRSI driven by a drift anisotropy, and thus bring in a new channel of self-confinement. This mechanism can potentially have many macroscopic consequences. For example in the case of CR bottlenecks (e.g, Skilling, 1971; Wiener et al., 2017) where the $\mathcal{E}_{\rm cr}$ gradient is flattened and the CRSI is deactivated, the CRPAI may then dominate and govern CR momentum transfer to the gas, which may alter the properties of CR-driven wind. In situations such as a supernova remnant where the system undergoes rapid expansion, the CR pressure anisotropy could be efficiently driven and induce a high level of heating (Zweibel, 2020). Additionally, consider a magnetic field with a focusing structure and suppose there is no MHD wave scattering at all, then CR flux should still be affected by the $\Delta P_{\rm cr}$ in the advection term $\nabla\cdot\mathbf{P}_{\rm cr}$ due to adiabatic invariance.
@@
Expanding CR distributions $f$ by Legendre polynomials is canonical in the study about CR transport because of their natural geometric characteristics in pitch angle (e.g., Klimas and Sandri, 1971; Earl, 1973; Webb, 1987; Zank et al., 2000; Snodin et al., 2006; Litvinenko and Noble, 2013; Rodrigues et al., 2019; Thomas and Pfrommer, 2019; Zweibel, 2020). Mathematically, a rigorous expansion certainly must involve a complete set of Legendre polynomials as basis functions. However, from the definitions in Equation 11, Equation 14, Equation 18 and the relation ${\rm tr}(\mathbf{P}_{\rm cr})=\mathcal{E}_{\rm cr}$, we see Legendre polynomials up to the second order are already sufficient to incorporate $\Delta P_{\rm cr}$. We thus truncate at second order. Given the orthogonality of Legendre polynomials, each $f_{n}(p,t)\ (n=0,1,2)$ component separately corresponds to a fluid moment:
@@
With these conventions, we now take the $\mathcal{E}_{\rm cr}$ and $F_{\rm cr,\parallel}/c^{2}$ moments on top of the kinetic Equation 3, yielding left hand sides (LHSs) as:
@@ II.3 CR Scattering
The values of these terms depend on the intensities of the corresponding resonant waves (Schlickeiser, 1989), which are determined by specific physical processes. However, even without a specified wave-driving mechanism, from the resonant condition $k\mu v\approx\pm\Omega_{\rm cr}/\gamma$ we can see forward/backward traveling particles $(\mu>0/\mu<0)$ only resonantly interact with right-hand/left-hand polarized waves $(+/-\ {\rm in\ front\ of\ }\Omega_{\rm cr})$, such that:
@@
The exact value of the factor $\alpha$ can be calibrated from MHD-PIC simulations (e.g., Bai, 2022; Sun and Bai, 2023).
@@
$+5\alpha c[(\sigma_{+}^{L}-\sigma_{+}^{R})w_{+,\parallel}+(\sigma_{-}^{L}-\sigma_{-}^{R})w_{-,\parallel}]\Delta P_{\rm cr}$
@@
$+5\alpha c\left(\sigma_{+}^{L}-\sigma_{+}^{R}+\sigma_{-}^{L}-\sigma_{-}^{R}\right)\Delta P_{\rm cr}$
@@
II.4 CR Hydrodynamic Equations in Lab Frame
@@
As CRs are relativistic particles, CR hydrodynamics shares similarities to RHD particularly in frame transformation relations. Thomas and Pfrommer (2019) showed the results for CRs by combining $\mathcal{E}_{\rm cr}$, $\vec{F}_{\rm cr}$ and $\mathbf{P}_{\rm cr}$ as a rank-2 energy momentum tensor and then applying a Lorentz transformation. We slightly extended their results to the case with an anisotropic $\mathbf{P}_{\rm cr}$, and list the relations below. Our calculation and final results are in direct analogy with the relations in RHD (e.g., Mihalas and Mihalas, 1984; Castor, 2007; Jiang et al., 2012):
@@
One issue associated with transforming $\mathbf{P}_{\rm cr}$ to the lab frame is the difficulty of determining $\Delta P_{\rm cr}$ in the lab frame. In fact it makes even no sense to define $\Delta P_{\rm cr}$ in the lab frame because the CR distribution $f$ is no longer symmetric about the gyro phase. However, quasi-linear theory (QLT) suggests that the anisotropic component in the ansatz Equation 19 will approximately satisfy $|f_{1}/f_{0}|\sim v_{A}/c$ and $|f_{2}/f_{0}|\sim v_{A}/c$ (Sun et al., 2024), which by equations Equation 20-Equation 22 implies $|\vec{F}_{\rm cr}|\sim v_{A}\mathcal{E}_{\rm cr}$ and $|\Delta P_{\rm cr}|\sim v_{A}\mathcal{E}_{\rm cr}/c$ (see also Section Section II.5). Consequently, the additional term arising from $\mathbf{P}_{\rm cr}$ transformation is $\sim\mathcal{O}(v_{A}u_{\perp}/c^{2})$, which is generally significantly smaller than $\Delta P_{\rm cr}$. We can thus consider $\mathbf{P}_{\rm cr,lab}\approx\mathbf{P}_{\rm cr}$, and regard $\Delta P_{\rm cr}$ somewhat as a scalar. These considerations imply that the closure for $\mathbf{P}_{\rm cr,lab}$ becomes
@@
Thomas and Pfrommer (2019) suggested this Lorentz force should counteract the perpendicular CR inertia and establish a dynamical equilibrium $\vec{j}_{\rm cr}\times\vec{B}/c=\nabla_{\perp}P_{\rm cr}$ on a kinetic time scale, which keeps the $\vec{F}_{\rm cr}$ in the comoving frame parallel to the magnetic field. Given such a consideration and the fact that Equation 24 is obviously the parallel component of $c^{-2}\partial\vec{F}_{\rm cr,lab}/\partial t+\nabla\cdot\mathbf{P}_{\rm cr,lab}$, we expect the additional perpendicular component to balance the Lorentz force, and thus write down the full CR flux equation as:
@@
$+\left[(\sigma_{+}^{L}-\sigma_{+}^{R})(u_{\parallel}+v_{A})+(\sigma_{-}^{L}-\sigma_{-}^{R})(u_{\parallel}-v_{A})\right]c\Delta P_{\rm cr}$
@@ II.5 Subgrid Physics
    A large-scale CR pressure gradient $\nabla P_{\rm cr}$ drives CR bulk motion (streaming), which triggers CRSI once this bulk motion exceeds the Alfvén speed. The CRSI excites two branches of the Alfvén waves propagating down the direction of $\nabla P_{\rm cr}$ projected along the magnetic field (Kulsrud and Pearce, 1969; Wentzel, 1974).
    Assuming a steady balance between wave damping and CRSI-driven growth, the scattering coefficients have been shown to be (Kulsrud, 2004; Bai et al., 2019; Bai, 2022; Armillotta et al., 2021):
@@
    corresponding to the case $\nabla P_{\rm cr}\cdot\vec{B}<0$ where Alfvén waves along the magnetic field are driven. For $\nabla P_{\rm cr}\cdot\vec{B}>0$, the non-zero contribution instead lies in the $\sigma_{-}^{L,R}$ branch with the same magnitude. Here $\nu_{\rm damp}$ is the damping rate, and $\rho$ is the ionized gas density.
    With this prescription, the resulting equations are consistent with the current two-moment CR fluid formulation (e.g. Jiang and Oh, 2018; Thomas and Pfrommer, 2019) as expected in the situation solely subject to the CRSI. The above relation applies for either linear or nonlinear wave damping (e.g. ion-neutral or nonlinear Landau; (Kulsrud and Pearce, 1969)). For ion-neutral damping, appropriate in neutral gas, the resulting $\sigma_{\pm}^{L,R}$ is linear in $|\nabla P_{\rm cr}|$; for nonlinear Landau damping, appropriate in ionized gas, $\sigma_{\pm}^{L,R}$ varies as $\sqrt{|\nabla P_{\rm cr}|}$ (Armillotta et al., 2021).
@@
    The CRPAI is a less studied version of the CR gyro-resonant instability, which can be triggered when the level of CR pressure anisotropy $\Delta P_{\rm cr}/P_{\rm cr}\gtrsim v_{A}/c$ (Kulsrud, 2004; Zweibel, 2020). It triggers one branch of Alfvén wave along each propagating direction. In reality, this may occur when the bulk CRs experience a varying magnetic field, e.g., when CRs travel through magnetic fields with large spatial variations, or the background field undergoes shear motion. Recent work (e.g., Sun et al., 2024) has isolated this effect using an expanding/compressing box setup. They derived the scattering coefficients based on QLT, and calibrated the results using MHD-PIC simulations. The result is:
@@
    This applies for $\Delta P_{\rm cr}<0$ (increasing field strength). If instead $\Delta P_{\rm cr}>0$, then the roles reverse: $\sigma_{+}^{L}=\sigma_{-}^{R}=0$, while $\sigma_{+}^{R}=\sigma_{-}^{L}$ takes the non-zero value given above.
@@
The other ingredient for user input is the value of CR pressure anisotropy $\Delta P_{\rm cr}$, which is again associated with the CRPAI. Similarly based on QLT, by balancing wave growth and (linear) damping, together with the growth of pressure anisotropy and quasi-linear diffusion, the steady state value of $\Delta P_{\rm cr}$ can be estimated as (Sun et al., 2024):
@@
The +/+/- signs correspond to a decreasing/increasing magnetic field strength that makes $P_{\rm cr,\perp}$ smaller/larger than $P_{\rm cr,\parallel}$.
We highlight two important points behind the aforementioned prescriptions. First, the scattering coefficients and CR pressure anisotropy are derived assuming steady-state balance between wave driving and damping. Typically, this balance can be established on an instability growth timescale $\sim(n_{\rm ion}/n_{\rm cr})\Omega_{\rm cr}^{-1}\sim 10^{3-4}\ \rm yr$, which is orders of magnitude shorter than the dynamic timescale of the macroscopic system. Therefore, it usually suffices to adopt such steady-state estimates. Second, we note that these prescriptions were calculated by assuming either the CRSI or CRPAI being the only instability at work. In reality, the CRPAI and CRSI are likely to act together, although the outcome is not additive, because each of them amplify (different) two out of four parallel-propagating Alfvén waves, while damping the other two. For present purposes, we only consider the case with pure CRSI or CRPAI, while leaving the more general case to future work.
III Numerical Implementation